• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 08:00:29 on October10,2010

    Entry number 334187

    keyword=Owl shift summary

    12:14 - MCC called. They will begin to set up Hall C, and we may loose beam for 30-60 minutes.

    12:31 - Since there's no beam, Alex wants to take a quick controlled access to remove a board from a crate.

    1:05 - Hall back to beam permit.

    1:30 - MCC can now deliver beam again.

    1:56 - MCC has an ARC power supply problem. Beam will be down for at least 1 hour.

    4:30 - MCC almost ready to deliver beam. Go to empty target for beam recovery.

    5:05 - MCC is actually ready to deliver beam. Start with beam recovery.

    5:20 - ARC power supply is down again. Luckily, the experts are still here from the last down, so it should be a quick fix...

    During the last couple of hours of the shift, there have been multiple quick left arm NMR alarms. It has always recoverd, so far...

    Ending the shift still with no beam.

    Optics runs:

    p0=3.194GeV/c (p0=-4%):
    Run 7099, 296k events
    Run 7106, only first 20 minutes had beam